# OVSeg3R: Learn Open-vocabulary Instance Segmentation from 2D via 3D Reconstruction

### v1

@@ A.1 More Ablations
Ablation on 3D Reconstruction Quality Although 3D reconstruction methods have advanced considerably (Wang et al., 2024; Leroy et al., 2024; Murai et al., 2025; Wang et al., 2025), failures still occur due to errors in camera parameters or depth estimation. To assess their impact on the model, we keep the failed reconstructions and conduct an ablation study. As shown in Table 4, filtering out failed reconstructions yields a clear performance gain, demonstrating the potential of OVSeg3R. With continued advances in 3D reconstruction, as accuracy improves and failure rates decrease, the contribution of OVSeg3R in the 3D scene understanding will be further enhanced.
Ablation on Input 2D Features Since our experiments are based on SegDINO3D-VL, an extension of SegDINO3D (Qu et al., 2025), which proposes to leverage well-trained 2D features to help data-hungry 3D models in understanding 3D scenes. Thus, as in SegDINO3D, input 2D image- and object-level features are important. By default, we use DINO-X (Ren et al., 2024) in universal mode [^2] In the universal mode of DINO-X, users do not need to provide text prompts specifying target classes, DINO-X automatically detects all objects. to provide features. For comparison, we also extract features using DINO-X in the regular mode, restricting text prompts to the 20 ScanNet classes. As shown in Table 4, limiting the 2D model’s attention to these 20 classes prevents it from providing sufficient information for open-vocabulary recognition. While this has little negative impact on base classes (-1.2 mAP), it leads to a substantial performance drop on novel classes (-8.3 mAP).

### v2

@@ 6 Conclusion
7 Acknowledgement
This work is partially supported by Meituan Academy of Robotics Shenzhen.
@@ A.1 More Ablations
Ablation on 3D Reconstruction Quality Although 3D reconstruction methods have advanced considerably (dust3r; leroy2024grounding; mast3rslam; vggt), failures still occur due to errors in camera parameters or depth estimation. To assess their impact on the model, we keep the failed reconstructions and conduct an ablation study. As shown in Table 4, filtering out failed reconstructions yields a clear performance gain, demonstrating the potential of OVSeg3R. With continued advances in 3D reconstruction, as accuracy improves and failure rates decrease, the contribution of OVSeg3R in the 3D scene understanding will be further enhanced.
Ablation on Input 2D Features Since our experiments are based on SegDINO3D-VL, an extension of SegDINO3D (segdino3d), which proposes to leverage well-trained 2D features to help data-hungry 3D models in understanding 3D scenes. Thus, as in SegDINO3D, input 2D image- and object-level features are important. By default, we use DINO-X (DINOX) in universal mode [^2] In the universal mode of DINO-X, users do not need to provide text prompts specifying target classes, DINO-X automatically detects all objects. to provide features. For comparison, we also extract features using DINO-X in the regular mode, restricting text prompts to the 20 ScanNet classes. As shown in Table 4, limiting the 2D model’s attention to these 20 classes prevents it from providing sufficient information for open-vocabulary recognition. While this has little negative impact on base classes (-1.2 mAP), it leads to a substantial performance drop on novel classes (-8.3 mAP).